News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 13 September 2013 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

Chandra passed through the nineteenth, twentieth, and twenty-first (and final) eclipses of the season on Sep 6, 9, and 12 respectively, with nominal power and thermal performance. Real-time procedures were executed on Sep 12 to disable SCS-29 and to dump and clear the EPS glitch counters as post-eclipse season actions.

A real-time procedure was executed on Sep 6 to turn-off Fine Sun-Sensor A (FSS-A) and Sensor Processing Electronics B (SPE-B). FSS-A and SPE-B had been left powered on since the swap to using FSS-B for routine operations on May 10 in order to allow for additional monitoring of FSS-A performance. At this time adequate data have been collected to be confident of the use of FSS-A for safe-mode operation.

A real-time procedure was used to monitor the momentum unload that was executed from the on-board command loads on Sep 6; the unload performance was nominal.

The schedule of targets for the next week is shown below and includes an observation of GRB130831A, which was accepted as a Director's Discretionary Time TOO on Sep 9, observations of SN 2012ca and SN 2013ej, which were accepted as a Targets of Opportunity on Aug 14 and Sep 4 respectively, and an observation of SN 2012ca coordinated with Spitzer.



------------------------------------------
        PSRJ0633+1746          ACIS-I       Sep 16
        ArLac (5 obs)          HRC-I
        ArLac (4 obs)          HRC-S
        ArLac (4 obs)          HRC-I
        ArLac (4 obs)          HRC-S
        ArLac (4 obs)          HRC-I
        GRB130831A             ACIS-S       Sep 17
        Radiation Belts
        SN2012ca               ACIS-S
        ArLac (5 obs)          HRC-S
        ArLac (4 obs)          HRC-I        Sep 18
        ArLac (4 obs)          HRC-S
        ArLac (4 obs)          HRC-I
        ArLac (4 obs)          HRC-S
        G121.11+57.01          ACIS-I
        TCha                   ACIS-S/HETG
        G311.5-0.3             ACIS-I       Sep 19
        TCha                   ACIS-S/HETG
        Radiation Belts                     Sep 20
        SgrA*                  ACIS-S
        PSRJ0633+1746          ACIS-I
        SN2013ej               ACIS-S       Sep 21
        HE0435-1223            ACIS-S
        HAE63                  ACIS-S       Sep 22
        PSZ226                 ACIS-I

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports